### v2

Abstract
Feature selection is critical for improving the performance and interpretability of machine learning models, particularly in high-dimensional spaces where complex feature interactions can reduce accuracy and increase computational demands. Existing approaches often rely on static feature subsets or manual intervention, limiting adaptability and scalability. However, dynamic, per-instance feature selection methods and model-specific interpretability in reinforcement learning remain underexplored. This study proposes a human-in-the-loop (HITL) feature selection framework integrated into a Double Deep Q-Network (DDQN) using a Kolmogorov-Arnold Network (KAN). Our novel approach leverages simulated human feedback and stochastic distribution-based sampling, specifically Beta, to iteratively refine feature subsets per data instance, improving flexibility in feature selection. The KAN-DDQN achieved notable test accuracies of 93% on MNIST and 83% on FashionMNIST, outperforming conventional MLP-DDQN models by up to 9%. The KAN-based model provided high interpretability via symbolic representation while using 4 times fewer neurons in the hidden layer than MLPs did. Comparatively, the models without feature selection achieved test accuracies of only 58% on MNIST and 64% on FashionMNIST, highlighting significant gains with our framework. We further validate scalability on CIFAR-10 and CIFAR-100, achieving up to 30% relative macro F1 improvement on MNIST and 5% on CIFAR-10, while reducing calibration error by 25%. Complexity analysis confirms real-time feasibility with latency below 1 ms and parameter counts under 0.02M. Pruning and visualization further enhanced model transparency by elucidating decision pathways. These findings present a scalable, interpretable solution for feature selection that is suitable for applications requiring real-time, adaptive decision-making with minimal human oversight. The code is available at https://github.com/Abrar2652/HITL-FS.
{IEEEkeywords}
Human-in-the-Loop, Feature Selection, Kolmogorov-Arnold Network, Reinforcement Learning, Double Deep Q-Network.
1 Introduction
[…]
Feature selection is essential for building efficient, accurate, and robust machine learning models [6, 2]. While models ideally should automatically identify the most predictive features, a high-dimensional input space can significantly hinder performance, often requiring large volumes of data to effectively learn the complex relationships between features. This phenomenon, known as the “curse of dimensionality,” increases computation time and resource use. Consequently, selecting a smaller subset of relevant features improves performance and makes the model more cost-effective.
One common solution is to incorporate expert knowledge to determine the most useful features; however, this process is costly, time-consuming, and highly manual. Additionally, experts with deep domain knowledge are often not involved in the actual design and development of the model. Automatic feature selection methods offer an alternative, ranking features by their relevance or importance; however, both manual and automatic approaches typically yield a single subset of features for the entire dataset, which may not capture variability across individual observations. When training data are sparse relative to the feature space, using a single subset can limit the model’s ability to generalize effectively across all instances [3].
To address the challenge of per-example feature selection, we propose a reinforcement learning (RL) framework that leverages simulated feedback to replicate human feature selection. Our approach employs a Double Deep Q-Network (DDQN) setup with a Kolmogorov-Arnold Network (KAN) [8, 5] as both the Q-network and the target network. This model-specific, interpretable KAN-based RL structure aims to refine feature selection on a per-example basis iteratively. In this setup, simulated feedback, rather than direct human input, is a proxy for expert annotation [1, 4]. The feedback signal highlights the most relevant features for each data example, enabling the model to prioritize these features during training. Unlike prior methods [3], our model explores distributions beyond Bernoulli and organizes convolutional and pooling layers to ensure feature maps match the input and simulated feedback shapes. RL then optimizes a policy to select a unique subset of features per observation. By minimizing the classifier’s prediction loss and the discrepancy between simulated feedback and the model-selected features, this policy yields feature subsets that improve the interpretability and performance of the final predictions. Since predictions are based only on the selected feature subsets, this method offers interpretable, case-specific insights into the model’s output. Using simulated feedback further enables the model to reflect causal structures likely to be relevant in practical applications. We validate our methodology through rigorous experimentation on benchmark datasets, showing the efficiency of our approach in improving model accuracy while maintaining computational feasibility. We aim to establish best practices for integrating human feedback into the feature selection process by investigating the influence of various hyperparameters, stochastic distributions, and the absence of feature selection.
In many deployed classification settings, such as safety triage in vision, early exits in mobile perception, and clinician-in-the-loop screening, models must meet tight memory/latency budgets while preserving case-wise interpretability. Static, dataset-level feature subsets often underperform when intra-class variability is high; conversely, per-instance selection can reduce redundant computation and expose the causal drivers of individual decisions. Our goal is to operationalize instance-wise feature selection under realistic constraints: (i) a compact head suitable for resource-limited environments, (ii) a stochastic, differentiable gating mechanism that trades sparsity for accuracy, and (iii) a reinforcement-learning controller whose policy remains interpretable. This motivates our HITL-DDQN framework, which features a KAN or MLP head and a stochastic gate aligned with simulated feedback, thereby delivering actionable sparsity–accuracy operating points and model-specific interpretability.
The key components of our contribution include the following:
  1. 1.
    We introduce a novel approach incorporating simulated feedback via Gaussian heatmaps and stochastic, distribution-based sampling to refine feature subsets on a per-example basis, thereby enhancing model interpretability and performance.
  2. 2.
    By incorporating KAN into the DDQN architecture for both the Q-network and target network, we achieve significantly better performance than traditional MLP-based DDQN across all test cases. This approach uses a hidden layer with four times fewer neurons than MLP while offering model-specific interpretability.
  3. 3.
    Our research presents a simulated feedback mechanism that generates feature relevance feedback without the need for human annotators, facilitating a scalable training process that reflects the causal relationships typically identified by human experts.
The following sections provide an overview of our work: Section 2 reviews the relevant background; Section 3 outlines the complete methodology of our proposed framework; Section 4 details the experimental design; Section 5 discusses the results and their interpretations; and finally, Section 6 concludes the research while outlining potential future directions.
2 Background
2.1 Human-in-the-Loop Feature Selection Using RL
Feature selection is critical in developing machine learning models but is often executed through data-driven methods that overlook insights from human designers [2, 9]. We introduce a HITL [16] framework that integrates simulated feedback to identify the most relevant variables for specific tasks, which can be modeled using DDQN-based [7] RL to facilitate per-example feature selection [13], enabling the model to minimize its loss function while emphasizing significant variables from a simulated human perspective. A major gap in RL is the limited model-specific interpretability, as most models operate as black-box systems, making it challenging to understand their decision-making processes [14]. Our KAN agent enhances the interpretability of DDQN-based RL by providing symbolic representations of learned policies. Our methodology employs variable elimination techniques, focusing on selecting subsets of features rather than using embedded methods. This approach optimizes feature selection and learning processes concurrently via gradient descent, distinguishing itself by selecting different subsets for each observation. This enhances interpretability, as the chosen variables represent the “causes” driving the model’s predictions. While previous work [6, 16] has explored per-example feature selection via traditional filter methods based on mutual information, our framework goes beyond incorporating simulated human feedback for a more dynamic approach. Traditional feature selection techniques include filter methods, which select top-ranked features based on criteria like mutual information; wrapper methods, which evaluate subsets of features by retraining models for each subset; and embedded methods, which attempt to select features during the model training process. Unlike these approaches, which can be sequential and computationally costly, our model generates candidate subsets in a single step, guided by simulated feedback, thereby avoiding arbitrary stopping criteria and allowing for real-time, per-example selection in complex datasets. Our work draws inspiration from the probabilistic knowledge elicitation [1, 15] focused on querying users for global feature relevance. However, our model goes beyond the focus on Bernoulli distributions [3] by exploring various distributions, including beta, Gaussian, and Gumbel-Softmax, enhancing the flexibility and effectiveness of feature selection. This breadth allows us to capture a wider range of relationships within the data compared with previous approaches.
2.2 Kolmogorov-Arnold Networks (KANs)
KANs [11] utilize the Kolmogorov-Arnold representation theorem, which states that any multivariate continuous function $f:[0,1]^{n}\to\mathbb{R}$ can be expressed as a finite composition of univariate functions and additions. Mathematically, this is formulated as follows:
$f(x)=\sum_{q=1}^{2^{n+1}}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right)$ (1)
where $\phi_{q,p}:[0,1]\to\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are continuous functions. In KANs, the traditional weight parameters are substituted with learnable one-dimensional B-spline functions $\phi$. The output of a KAN layer with $n_{in}$ inputs and $n_{out}$ outputs is given by:
$x_{l+1,j}=\sum_{i=1}^{n_{l}}\phi_{l,j,i}(x_{l,i})$ (2)
where $\phi_{l,j,i}$ connects the $i$-th neuron in layer $l$ to the $j$-th neuron in layer $l+1$. The backpropagation algorithm computes gradients with respect to the spline coefficients $c_{i}$ to minimize the loss $L$ via gradient descent:
$\frac{\partial L}{\partial c_{i}}=\sum_{j=1}^{n_{out}}\frac{\partial L}{\partial x_{l+1,j}}\frac{\partial x_{l+1,j}}{\partial c_{i}}$ (3)
where $\frac{\partial x_{l+1,j}}{\partial c_{i}}$ reflects the derivative of the spline function concerning its coefficients. KANs define spline functions $\phi$ over a discretized grid, which specifies where the functions are evaluated, thus influencing the approximation resolution. The order parameter indicates the degree of the B-splines: an order of 1 denotes linear splines, while higher orders yield more complex shapes. The grid and order parameters determine KANs’ capacity to model intricate functional relationships; finer grids with higher-order splines allow for precise approximations but require more computational resources.
The primary difference between KAN and MLP is that KANs implement learnable activation functions along edges, whereas MLPs implement fixed activation functions at nodes. KANs represent weights as splines, improving their ability to approximate complex functions with fewer parameters. KANs and MLPs can be extended to multiple layers, supporting deep architectures. We introduce KANs here to frame their conceptual differences from MLPs; the practical instantiation inside our DDQN head (width, grid, regularization) is given later in Sec. 3-3.3 to avoid duplication and to keep implementation choices distinct from background theory.
Figure 1: System overview. The feature-selection head (FSNet) maps an input image to per-feature probabilities via a differentiable stochastic gate aligned to simulated feedback. The selected features are fed to a DDQN head (KAN or MLP). Replay-based training updates the Q-network, with periodic target sync. The gate induces instance-wise sparsity and exposes feature-level rationale, while the KAN head provides model-specific interpretability.
3 Methodology
Input: Input shape input_shapeinput\_shape, Number of filters num_filtersnum\_filters, Number of convolutional layers num_conv_layersnum\_conv\_layers, Hidden layer dimension hidden_dimhidden\_dim, Distribution distributiondistribution, Temperature parameter $\tau$
Result: Processed feature map, feedback cost, feature probabilities
1
21ex
3 Function FSNet(input_shapeinput\_shape, num_filtersnum\_filters, num_conv_layersnum\_conv\_layers, hidden_dimhidden\_dim, distributiondistribution, $\tau$):
4  Initialize convolutional layer parameters: $channels\leftarrow$ []; for $i\leftarrow 1$ to num_conv_layersnum\_conv\_layers do
5     $num\_filters\leftarrow num\_filters\times 2$;
6     Append num_filtersnum\_filters to channelschannels;
7
8  end for
9 Build feature extractor with num_conv_layersnum\_conv\_layers Conv2D, ReLU, and MaxPool layers;
10  Initialize fully connected layer $fc1$ with input size n_featuresn\_features and hidden size hidden_dimhidden\_dim;
11  Initialize output layer $fc2$ with input size hidden_dimhidden\_dim and output size n_featuresn\_features;
12  Initialize activation functions: relurelu, sigmoidsigmoid, and softmaxsoftmax;
13
14
15
161ex
17 Function forward($x$, feedbackfeedback, epochepoch):
18  Reshape $x$ to match input_shapeinput\_shape;
19  $features\leftarrow$ Pass $x$ through feature extractor;
20  $x\leftarrow relu(fc1(features))$;
21  Compute $logits\leftarrow relu(fc2(x))$;
22  Compute feature selection probabilities $probs\leftarrow sigmoid(logits)$;
23
24 if feedbackfeedback is None then
25     $feedback\leftarrow probs$;
26
27  end if
28
29 Sample sample_probssample\_probs from distributiondistribution using logitslogits as parameters;
30  Adjust temperature parameter if Gumbel-Softmax is used: $tau\leftarrow\tau\times\tau\_decay^{epoch}$;
31
32 Calculate feedback cost $feedback\_cost\leftarrow MSE\_Loss(probs,feedback)$;
33  $selected\_feature\leftarrow features\times sample\_probs$;
34
35 return selected_featureselected\_feature, feedback_costfeedback\_cost, probsprobs;
36
37 return
Algorithm 1 FSNet Model Initialization and Forward Pass
3.1 Human-in-the-Loop Feature Selection Process
The feature selection network (FSNet) (see Algorithm 1 and Fig. 1) is designed to determine the relevance of each input feature for a given task [13]. The architecture combines convolutional layers for feature extraction with various probabilistic distributions to model and sample feature importance. Here, we detail each model component, focusing on architectural elements, distribution-based sampling, and feedback alignment.
3.1.1 Convolutional Feature Extraction Layers
The first stage of FSNet involves convolutional layers that transform the input tensor $x$ into feature representations that capture important patterns across spatial dimensions. Suppose $x\in\mathbb{R}^{d\times h\times w}$ represents an input with $d$ channels, height $h$, and width $w$. The convolutional feature extractor uses $L$ layers of convolution, ReLU activation, and max-pooling to refine the spatial features of the input progressively. For layer $i$, with filter count $f_{i}$, the transformation can be expressed as:
$F^{(i)}=\text{MaxPool2d}(\text{ReLU}(\text{Conv2d}(F^{(i-1)})))$ (4)
where $F^{(i)}\in\mathbb{R}^{f_{i}\times\frac{h}{2^{i}}\times\frac{w}{2^{i}}}$ and $F^{(0)}=x$. Each convolutional layer uses a $3\times 3$ kernel and padding of 1 to maintain spatial dimensions, while max-pooling layers with a kernel size of 2 reduce the height and width by half at each step.
3.1.2 Fully Connected Layers for Feature Probability Mapping
After convolutional layers, the resulting feature map is flattened and processed by fully connected layers to output feature relevance scores (logits). Let $F\in\mathbb{R}^{n}$ represent the flattened feature vector, where $n$ is the number of features output by the convolutional layers. This vector is then passed through a sequence of fully connected (FC) layers, ReLU activation, and dropout:
$h=\text{ReLU}(W_{\text{fc1}}\cdot F+b_{\text{fc1}})$ (5)
where $h\in\mathbb{R}^{128}$ (if 128 hidden units are used), $W_{\text{fc1}}\in\mathbb{R}^{128\times n}$ is the weight matrix, and $b_{\text{fc1}}\in\mathbb{R}^{128}$ is the bias vector. The ReLU introduces non-linearity, while dropout (with a rate of 0.25) mitigates overfitting. Finally, the output layer applies a sigmoid function to yield probability values for each feature:
$\hat{q}=\sigma(W_{\text{fc2}}\cdot h+b_{\text{fc2}})$ (6)
where $\hat{q}=(\hat{q}_{1},\hat{q}_{2},\ldots,\hat{q}_{d})$ represents the probability score for each feature. The sigmoid activation ensures that $\hat{q}_{j}\in[0,1]$ for all $j$, making $\hat{q}$ interpretable as a relevance score or probability vector and $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function.
3.1.3 Distribution-Based Sampling for Probabilistic Feature Selection
[…]
$\pi(a\mid\hat{q})=\prod_{j=1}^{d}P(a_{j}\mid\hat{q}_{j})$ (7)
where $P(a_{j}\mid\hat{q}_{j})$ denotes the probability of selecting (or discarding) feature $j$ conditioned on its relevance score $\hat{q}_{j}$. This probabilistic formulation introduces flexibility by allowing the network to dynamically control the inclusion probability of each feature through its learned relevance estimate.
To parameterize the feature selection probabilities, we employ the Beta distribution, a continuous distribution bounded within $[0,1]$, which is well-suited for modeling probabilities, as shown in Table 1. Each feature’s selection probability is sampled as $a_{j}\sim\text{Beta}(\alpha_{j},\beta_{j})$, where the shape parameters are computed as:
$\alpha_{j}=\beta_{j}=\text{softplus}(\hat{q}_{j})+1$ (8)
Here, $\text{softplus}(x)=\ln(1+e^{x})$ ensures positivity and numerical stability for the shape parameters. The symmetry $\alpha_{j}=\beta_{j}$ allows the distribution to smoothly vary from uniform to highly skewed based on the learned relevance score $\hat{q}_{j}$. When $\hat{q}_{j}$ is large, the Beta distribution skews toward 1 (favoring feature retention), and when small, it skews toward 0 (favoring feature exclusion). This approach provides a highly flexible and differentiable mechanism for stochastic feature selection.
The sampled feature mask $\mathbf{a}$ is then applied to the feature vector $\mathbf{f}$ as $\mathbf{f}_{\text{selected}}=\mathbf{f}\odot\mathbf{a}$, where $\odot$ denotes element-wise multiplication, yielding a filtered feature vector $\mathbf{f}_{\text{selected}}$ that selectively retains features based on their relevance.
3.1.4 Feedback Alignment with MSE Loss
FSNet uses mean squared error (MSE) to align the model’s feature relevance scores with human-provided feedback. Let $f\in\mathbb{R}^{d}$ represent a feedback vector where each $f_{j}\in[0,1]$ indicates human-assessed importance for feature $j$. The feedback cost function $C_{f}$ is defined as:
$\mathcal{L}_{\text{feedback}}=C_{f}(x;\hat{q},f)=\mathbb{E}\left[\|f-\hat{q}\|^{2}\right]=\frac{1}{d}\sum_{j=1}^{d}(f_{j}-\hat{q}_{j})^{2}$ (9)
This loss function penalizes discrepancies between the model’s probability vector $\hat{q}$ and the feedback $f$. Minimizing $C_{f}$ encourages FSNet to produce relevance scores that align with human intuition, resulting in a more interpretable feature selection.
Input: Configuration config, containing training hyperparameters
Output: Trained Q-network, performance metrics
1 Initialize environment $\mathcal{E}$, replay buffer $\mathcal{B}$ with capacity buffer_size;
2 Define $Q_{\theta}$, the Q-network, and $Q_{\theta^{\prime}}$, the target network;
3 Initialize optimizer with learning_rate and weight_decay;
4 Define the learning rate scheduler with decay factor $\gamma$;
5
6for each epoch $t=1,2,\ldots,\textit{n\_epochs}$ do
7  Initialize training metrics: running loss, correct predictions, feedback cost;
8  for each batch $b$ in environment $\mathcal{E}$ do
9     Obtain current state $s$, label $y$, and feedback from environment;
10     if feature_selection is True then
11        Apply feature selection using agent, compute feedback cost and probabilities;
12        Store feedback cost and probabilities;
13        Update $s\leftarrow\text{processed state}$;
14
15     end if
16    else
17        Use raw state;
18
19     end if
20
21    Choose action $a$ using $\epsilon$-greedy on $Q_{\theta}$ or random action if warm-up;
22     Compute reward based on $a$ and $y$;
23     Obtain next state $s^{\prime}$, apply feature selection if enabled;
24     Store $(s,a,s^{\prime},r)$ in $\mathcal{B}$;
25
26    if buffer $\mathcal{B}$ is ready (size $\textit{batch\_size}$) then
27        Sample a batch from $\mathcal{B}$;
28        Compute $Q$-learning target $y=r+\gamma\max_{a^{\prime}}Q_{\theta^{\prime}}(s^{\prime},a^{\prime})$;
29        Compute current estimate $Q_{\theta}(s,a)$;
30        Compute combined loss using $\mathcal{L}_{\text{SmoothL1}}$ with regularization;
31        if feature_selection is True then
32           Add feedback cost to loss;
33
34        end if
35       Backpropagate loss and update $Q_{\theta}$;
36
37     end if
38
39  end for
40 if $t\mod 25=0$ and method is ”KAN” and $t<\frac{\textit{n\_epochs}}{2}$ then
41     Update grid of $Q_{\theta}$ and $Q_{\theta^{\prime}}$ with samples from $\mathcal{B}$;
42
43  end if
44
45 if $t\mod\textit{target\_update}=0$ then
46     Update target $Q_{\theta^{\prime}}\leftarrow Q_{\theta}$;
47
48  end if
49 Adjust learning rate with scheduler;
50
51 end for
52Return $Q_{\theta}$ and collected metrics (accuracy, loss, feedback cost history);
Algorithm 2 Training Procedure for DDQN Model
3.2 Double Deep Q-Network Architecture
The DDQN architecture in our approach extends the traditional Q-learning framework by employing two neural networks: the primary $Q$-network, denoted as $Q_{\theta}$, and a target network, denoted as $Q_{\theta^{\prime}}$. The primary $Q$-network learns the action-value function $Q(s,a;\theta)$, estimating the expected cumulative reward for selecting action $a$ in a given state $s$. To stabilize training and mitigate the overestimation bias commonly observed in standard Q-learning, we use the target network $Q_{\theta^{\prime}}$, updated less frequently than the primary network.
During training, the parameters of $Q_{\theta}$ are updated via gradient descent, while the parameters of $Q_{\theta^{\prime}}$ are periodically synchronized with $Q_{\theta}$ to avoid target instability. To update $Q(s,a;\theta)$ toward the target (Equation 11), we leverage a replay buffer $\mathcal{B}$, which stores experience tuples $(s,a,r,s^{\prime})$. Sampling mini-batches from $\mathcal{B}$ helps reduce correlations between consecutive experiences, improving stability and allowing for independent updates to Q-values.
3.3 KAN
We now instantiate KAN as the DDQN head, following the background in Sec. 2-2.2.In our approach, the KAN-based architecture is structured with widths of [64, 8, 10] for the input, hidden, and output layers. Both the Q-network and target network use this compact configuration to capture complex feature interactions efficiently. Each KAN layer combines a spline function and a residual basis function, formulated as:
$\phi(x)=w_{b}b(x)+w_{s}spline(x)$ (10)
where $b(x)=silu(x)$ serves as a non-linear basis function, while $spline(x)=\sum_{i}c_{i}B_{i}(x)$ utilizes trainable B-splines for flexible approximation. With spline order $k=3$ over a specified grid, this layered structure dynamically adjusts to feature relevance throughout training. For optimization, $w_{b}$ and $w_{s}$ are initialized carefully: $w_{b}$ uses Xavier initialization, ensuring balanced layer activation scales, while $w_{s}$ is set to 1, with the initial spline function close to zero. The model’s spline grids adapt to changing input activations, extending KAN’s effective region for learning. The choice of KAN for the Q-network is theoretically motivated by its property of local plasticity, which mitigates catastrophic forgetting common in global MLP updates. We provide a detailed theoretical analysis of KAN approximation bias and gradient stability in Appendix 7.
3.4 MLP
The MLP-based architecture provides a more traditional setup for the DDQN framework, using a straightforward, fully connected design for both the Q-network and the target network. This configuration comprises two linear layers: a 64-neuron initial layer that maps the input features to a hidden space defined by a 32-neuron network width, followed by a ReLU activation, and a 10-neuron final linear layer that produces Q-values. The MLP requires a larger hidden layer with 32 neurons to approximate the complex mappings found in datasets, making it parameter-heavy relative to the KAN-based DDQN.
3.5 Training Procedure: KAN-based and MLP-based DDQN
The training procedure for the DDQN is illustrated in Algorithm 2 and Fig. 1, which outlines the main steps taken during each epoch. The algorithm initializes the environment and Q-networks and iteratively processes batches of experiences from a replay buffer $\mathcal{B}$. Each experience consists of the current state $s$, action $a$, reward $r$, and the next state $s^{\prime}$.
We employ two training algorithms to optimize our Q-network, each tailored to different model architectures: (1) a KAN-based training procedure that utilizes regularization for interpretability and stability, and (2) a standard MLP-based training procedure. Each training procedure integrates feedback cost when feature selection is enabled, promoting a minimalistic and interpretable feature set.
3.5.1 Temporal-Difference Target Computation
Both KAN and MLP-based approaches utilize the temporal-difference (TD) target to stabilize Q-learning updates and reduce the discrepancy between estimated Q-values and TD targets. For a given experience tuple $(s,a,s^{\prime},r,d)$, where $s$ and $s^{\prime}$ represent the current and next states, $a$ is the action taken, $r$ is the reward, and $d$ indicates the termination flag, the TD target is computed as:
$TD_{target}=r+\gamma\,Q_{\theta^{\prime}}\left(s^{\prime},\arg\max_{a}Q_{\theta}(s^{\prime},a)\right)$ (11)
where $\gamma\in[0,1]$ is the discount factor for future rewards. Here, $Q_{\theta}(s^{\prime},a)$ determines the action $a$ that maximizes the Q-value for the next state $s^{\prime}$, and $Q_{\theta^{\prime}}$ provides the stable estimate for this chosen action. The target network $Q_{\theta^{\prime}}$, updated less frequently, offers a stable estimation for TD updates.
3.5.2 KAN-based Training with Regularization
In the KAN-based training procedure, we employ $SmoothL1$ loss, which has proven effective in mitigating the influence of large outliers in Q-value errors. The primary loss term $\mathcal{L}_{\text{TD}}$ is given by:
$\mathcal{L}_{\text{TD}}=L_{\delta}\left(Q_{\theta}(s,a)-TD_{target}\right)$ (12)
where $L_{\delta}$ represents the $SmoothL1$ loss:
$L_{\delta}(x)=\begin{cases}0.5x^{2}&\text{if }|x|<1\\ |x|-0.5&\text{otherwise}\end{cases}$ (13)
L1 and Entropy-Based Regularization: In KANs, the L1 norm is applied to encourage sparsity in the activation functions, which replaces traditional linear weights used in MLPs. The L1 norm of an activation function $\phi$ is defined as the average magnitude of its outputs over $N_{p}$ inputs:
$\|\phi\|_{1}\equiv\frac{1}{N_{p}}\sum_{s=1}^{N_{p}}\phi(x^{(s)})$ (14)
For a KAN layer $\Phi$ with $n_{in}$ inputs and $n_{out}$ outputs, we define the L1 norm of $\Phi$ as the sum of the L1 norms of all activation functions:
$|\Phi|_{1}\equiv\sum_{i=1}^{n_{in}}\sum_{j=1}^{n_{out}}\|\phi_{i,j}\|_{1}.$ (15)
Additionally, an entropy term is introduced to mitigate overconfidence in the predictions. The entropy $S(\Phi)$ for the KAN layer is given by:
$S(\Phi)\equiv-\sum_{i=1}^{n_{in}}\sum_{j=1}^{n_{out}}\frac{|\phi_{i,j}|_{1}}{|\Phi|_{1}}\log\left(\frac{|\phi_{i,j}|_{1}}{|\Phi|_{1}}\right)$ (16)
The combined regularization term is defined as:
$\mathcal{R}_{\text{L1+Entropy}}=\sum_{i=1}^{n}\left(\lambda_{\text{L1}}\|\textit{acts\_scale}_{i}\|_{1}-\lambda_{\text{entropy}}H(\textit{acts\_scale}_{i})\right)$ (17)
where $H(\textit{acts\_scale}_{i})=-\sum p\log(p)$ denotes entropy with $p$ being the normalized activation values, and $\lambda_{\text{L1}}$ and $\lambda_{\text{entropy}}$ are hyperparameters.
Spline-Based Regularization: This term regularizes the spline coefficients of the KAN activation functions, encouraging smooth transitions and sparsity in the feature space. Given the spline coefficient vector $\text{coef}_{i}$ of each activation function, we compute the following:
$\mathcal{R}_{\text{Spline}}=\sum_{i=1}^{n}\left(\lambda_{\textit{coef}}\|\textit{coef}_{i}\|_{1}+\lambda_{\textit{coefdiff}}\|\textit{diff}(\textit{coef}_{i})\|_{1}\right)$ (18)
where $\textit{diff}(\textit{coef}_{i})$ calculates the adjacent differences within $\textit{coef}_{i}$, enforcing smoothness in the function and preventing rapid oscillations in the learned coefficients. The combined regularization term $\mathcal{R}$ is added to the loss as follows:
$\mathcal{L}_{\text{KAN}}=\mathcal{L}_{\text{TD}}+\lambda\mathcal{R}=L_{\delta}\left(Q_{\theta}(s,a),TD_{target}\right)+\lambda\left(\mathcal{R}_{\text{L1+Entropy}}+\mathcal{R}_{\text{Spline}}\right)$ (19)
3.5.3 Feedback Cost Integration for Feature Selection
To minimize unnecessary feature dependencies, we incorporate a feedback cost $\mathcal{L}_{\text{feedback}}$ (Equation 9) into the total loss with an $\alpha$ = 0.5 when feature selection is enabled. This cost penalizes selected features based on their contribution to the model, defined as follows:
$\mathcal{L}_{\text{KAN, total}}=\mathcal{L}_{\text{KAN}}+\alpha\times\mathcal{L}_{\text{feedback}}$ (20)
3.5.4 MLP-based Training
For the MLP-based training, we use the same TD target computation and $SmoothL1$ loss but omit the KAN-specific regularization terms. Like KAN-based training, feature selection introduces an additional feedback cost term to this loss: $\mathcal{L}_{\text{MLP, total}}=L_{\delta}\left(Q_{\theta}(s,a),TD_{target}\right)+\alpha\times\mathcal{L}_{\text{feedback}}$.
3.5.5 Optimization and Gradient Clipping
To ensure stable training and mitigate the risk of exploding gradients, we apply in-place gradient clipping with a threshold of 100 to both KAN and MLP-based training. This process limits the magnitude of gradients, preventing extreme updates that could destabilize the learning process: $\textit{clip}\left(\frac{\partial\mathcal{L}}{\partial\theta},100\right)$. The gradients are then backpropagated, and the optimizer updates the network weights to minimize $\mathcal{L}_{\text{KAN, total}}$ in KAN and $\mathcal{L}_{\text{MLP, total}}$ in MLP model.
3.6 Integration of HITL Feedback in DDQN
Integrating HITL feedback into the DDQN architecture enables the model to iteratively improve its feature selection by incorporating human expertise. This process adjusts the feature selection probabilities $\hat{q}_{j}$ based on human feedback, leading to an improved and rationale-aware selection policy.
3.6.1 Feature Selection Adjustment Using HITL Feedback
For each feature $j$, HITL feedback provides a target value $f_{j}$, representing the importance of that feature according to human assessment. The DDQN uses this feedback to adjust the selected feature probabilities by minimizing the feedback cost $\mathcal{L}_{\text{feedback}}$, thereby aligning the DDQN’s policy with human expertise.
3.6.2 Action-Selection Policy with HITL Feedback
The DDQN framework selects actions via an $\epsilon$-greedy policy, which balances exploration and exploitation. At the start of training, $\epsilon$ is set to a high value, allowing the agent to explore actions and discover potentially valuable states randomly. Over time, $\epsilon$ gradually decays, encouraging the agent to exploit its learned policy by selecting actions that maximize the estimated Q-value from the primary $Q_{\theta}$ network:
$a=\begin{cases}\text{random action}&\text{with probability }\epsilon\\ \arg\max_{a^{\prime}}Q(s,a^{\prime};\theta)&\text{with probability }1-\epsilon\end{cases}$ (21)
This approach improves learning by promoting diverse experiences early on while progressively focusing on reliable, high-reward actions as training progresses. When HITL feedback is introduced, the DDQN updates its policy to emphasize features positively reinforced by feedback, using this guidance to improve feature selection iteratively and refine the policy across epochs. Therefore, this $\epsilon$-greedy mechanism prevents premature convergence to suboptimal policies, supporting more robust training for effective feature selection.
3.6.3 Iterative Learning Process
The training loop for DDQN, with HITL feedback, ensures that the Q-network $Q_{\theta}$ converges to an optimal feature selection policy. At each epoch, the Q-network’s weights are updated based on both the Q-learning target and the feedback cost. This iterative process continues as:
$\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{total}}$ (22)
where $\eta$ is the learning rate and $\alpha$ is a hyperparameter balancing the feedback cost and Q-network loss. By adjusting $\alpha$ over time, the DDQN increasingly aligns its feature selection with human insights, resulting in a more interpretable and effective feature set.
3.6.4 Practical Implications and Robustness
While our framework utilizes simulated Gaussian feedback to model ideal expert attention, we discuss the practical implications, robustness to noisy annotations, and sensitivity to the loss function in Appendix 8.
4 Experimental Design
4.1 Dataset Preparation
We used a benchmark environment to standardize data preparation, feedback simulation, and visualizations across four datasets: MNIST, FashionMNIST, CIFAR-10, and CIFAR-100. Each dataset was resized to a standardized input dimension of $8\times 8$ pixels, with normalization applied based on empirically computed means and standard deviations specific to each dataset to aid in training convergence. For MNIST, we used a mean ($\mu$) of 0.1307 and a standard deviation ($\sigma_{SD}$) of 0.3081; for FashionMNIST, both $\mu$ and $\sigma_{SD}$ were set to 0.5. For CIFAR-10, normalization employed channel-wise $\mu$ of $(0.4914,0.4822,0.4465)$ and $\sigma_{SD}$ of $(0.247,0.243,0.261)$, while CIFAR-100 used $(0.5071,0.4865,0.4409)$ and $(0.2673,0.2564,0.2761)$ respectively. This normalization process adjusts pixel intensity $\mathbf{I}$ via the formula:
$\mathbf{I}_{\text{norm}}=\frac{\mathbf{I}-\mu}{\sigma_{SD}}$ (23)
where $\mu$ and $\sigma_{SD}$ are dataset-specific values.
Batch processing was performed with a batch size of $B=128$, enabling efficient data handling and parallel processing. Once a batch reaches its end, the iterator is reset, ensuring a continuous data stream throughout training. This setup streamlined the flow of images into memory along with the generated feedback signals for each batch.
Figure 2: A sample of preprocessed ($8\times 8$) pixel MNIST (left) and FashionMNIST (right) images and their corresponding feedback maps ($\sigma=5.0$).
4.2 Simulated Feedback Mechanism
To simulate supervisory guidance, we designed a feedback mechanism that highlights the most salient regions within each image via a Gaussian heatmap. The process begins with normalizing the input image $I\in\mathbb{R}^{C\times H\times W}$ to a range of $[0,1]$, defined by the transformation $I_{\text{norm}}=\frac{I-\min(I)}{\max(I)-\min(I)}$. This normalization ensures that all pixel values are scaled appropriately for subsequent computations. Next, the center of mass $C$ of the image is identified, represented by the coordinates $(x_{c},y_{c})$ corresponding to the pixel with the maximum intensity, calculated using $C=\text{argmax}(I_{\text{norm}})$. Utilizing these coordinates, a Gaussian feedback mask $M(x,y)$ is generated, which is centered around the identified pixel. The equation defines the Gaussian mask:
$M(x,y)=\exp\left(-\frac{(x-x_{c})^{2}+(y-y_{c})^{2}}{2\eta^{2}}\right)$ (24)
where $\eta=5.0$ is a parameter that controls the spread of the Gaussian distribution, reflecting the degree of uncertainty in the feedback. This mask provides spatial feedback, where values decay radially from the center (see Figure 2). To ensure that the feedback mask is interpretable and usable within the model, it is normalized so that its maximum value equals 1: $M_{\text{norm}}=\frac{M}{\max(M)}$. The resulting normalized feedback mask effectively emphasizes regions within the image with higher intensity values, which aligns to guide the model’s attention toward salient features.
4.3 Resources Used
Our experiments were conducted via a setup powered by an Intel(R) Xeon(R) CPU with a clock speed of 2 GHz, 4 virtual CPU cores, and 16 GB of DDR4 RAM. The software environment included Python 3.10.12, with PyTorch for model development, PyKAN for KAN, Gymnasium for environment simulations, SciPy and NumPy for numerical computations, Matplotlib for data visualization, and SymPy for symbolic calculations.
4.4 Experimental Configurations
For the experiments conducted in this study, we employed two configurations for the MLP and KAN models, both set to a batch size of 128 and trained over 100 epochs. The configurations utilized a learning rate $1\times 10^{-3}$, a weight decay $1\times 10^{-4}$, and a discount factor ($\gamma$) of 0.99. The MLP configuration featured a width of 32 and an output size of 10 classes, while the KAN configuration used a width of 8 and a grid size of 3. Both models share a common input size of 64 features, a buffer size of 100,000, and a target update interval of 10. They incorporated warm-up episodes of 2 and utilized a beta distribution (see Table 1) for their stochastic feature selection processes, with an initial $\tau$ value set at 1.0. We strategically designed the architecture of our model by selecting the appropriate combination of convolutional layers and filters to ensure that the resulting selected feature map retains the same shape as the input and feedback size (64 for $8\times 8$ images). This choice is crucial, as it directly allows the simulated feedback (similar to input size) to correspond to the feature map dimensions, facilitating effective integration during the feature selection process. In this case, for 1, 2, and 3 convolutional layers, we must select 2, 4, and 8 filter sizes to maintain the desired selected feature shape. Detailed descriptions of the evaluation metrics (accuracy, calibration, sparsity, and complexity), training of baselines, ablation protocols, and visualization scope are provided in Appendices 9.1, 9.2, 9.3, and 9.4.
Figure 3: (a) Training accuracy per epoch comparison of KAN and MLP-based DDQN on MNIST (left) and FashionMNIST (right), and (b) confusion matrix of KAN-DDQN on MNIST (left) and FashionMNIST (right).
Figure 4: (a) Pruned KAN architecture after removing neurons with both low incoming and outgoing activation magnitudes, exposing a minimal functional skeleton; (b) feature importances extracted from auto-symbolic forms of learned splines, indicating how often input features appear in simplified expressions; (c) early training trajectory (first five steps), where the policy concentrates mass on discriminative inputs before spreading to additional features. Together, these link sparsity, symbolic structure, and policy behavior.
Figure 5: Activation functions for middle neurons in the KAN agent’s hidden layer on the MNIST dataset. Each subplot shows how a specific neuron transforms inputs via the spline activation, illustrating various activation behaviors across the layer.
Table 1: Performance of MLP-DDQN for different distributions on MNIST and FashionMNIST datasets. The $\mu$ and $\sigma_{SD}$ are computed over a 10-epoch moving window. Bold indicates the best performance.
Distribution MNIST FashionMNIST
Train Accuracy Test Accuracy Train Accuracy Test Accuracy
Bernoulli 0.57 $\pm$ 0.03 0.58 $\pm$ 0.00 0.55 $\pm$ 0.02 0.54 $\pm$ 0.00
Gumbel Softmax 0.15 $\pm$ 0.01 0.18 $\pm$ 0.00 0.10 $\pm$ 0.02 0.24 $\pm$ 0.00
Gaussian 0.22 $\pm$ 0.01 0.21 $\pm$ 0.00 0.22 $\pm$ 0.01 0.21 $\pm$ 0.01
Beta 0.79 $\pm$ 0.02 0.78 $\pm$ 0.00 0.68 $\pm$ 0.01 0.66 $\pm$ 0.00
Dirichlet 0.17 $\pm$ 0.02 0.22 $\pm$ 0.00 0.23 $\pm$ 0.02 0.24 $\pm$ 0.00
Multinomial 0.10 $\pm$ 0.02 0.12 $\pm$ 0.00 0.10 $\pm$ 0.01 0.12 $\pm$ 0.00
Cauchy 0.12 $\pm$ 0.01 0.12 $\pm$ 0.00 0.11 $\pm$ 0.01 0.12 $\pm$ 0.00
Laplace 0.16 $\pm$ 0.02 0.18 $\pm$ 0.00 0.21 $\pm$ 0.01 0.19 $\pm$ 0.00
Uniform 0.73 $\pm$ 0.02 0.75 $\pm$ 0.00 0.69 $\pm$ 0.03 0.65 $\pm$ 0.00
Table 2: Performance Metrics without Feature Selection on MNIST and FashionMNIST Datasets Using KAN and MLP Models.
Model Dataset No Feature Selection Training Time (hh:mm:ss)
maP maR maF1 Test Accuracy
KAN MNIST 0.57 0.57 0.56 0.58 $\pm$ 0.01 5:10:50
MLP MNIST 0.45 0.51 0.47 0.52 $\pm$ 0.00 10:41
KAN FashionMNIST 0.63 0.65 0.63 0.64 $\pm$ 0.00 5:10:32
MLP FashionMNIST 0.55 0.60 0.57 0.59 $\pm$ 0.00 10:24
Table 3: Impact of FSNet parameters on test accuracy using MNIST and FashionMNIST datasets with feature selection
Model Dataset # Conv Layers # Filters # Width Grid Size With Feature Selection Training Time (hh:mm:ss)
maP maR maF1 Train Accuracy Test Accuracy
KAN MNIST 1 2 8 3 0.92 0.92 0.92 0.93 $\pm$ 0.01 0.93 $\pm$ 0.00 4:35:26
MLP MNIST 1 2 32 - 0.84 0.84 0.84 0.85 $\pm$ 0.02 0.84 $\pm$ 0.00 8:42
KAN FashionMNIST 1 2 8 3 0.83 0.83 0.83 0.85 $\pm$ 0.01 0.83 $\pm$ 0.00 4:25:41
MLP FashionMNIST 1 2 32 - 0.72 0.73 0.72 0.73 $\pm$ 0.01 0.74 $\pm$ 0.00 8:17
KAN MNIST 2 4 8 3 0.73 0.73 0.73 0.75 $\pm$ 0.02 0.73 $\pm$ 0.00 5:50:29
MLP MNIST 2 4 32 - 0.72 0.72 0.72 0.75 $\pm$ 0.02 0.72 $\pm$ 0.00 10:32
KAN FashionMNIST 2 4 8 3 0.69 0.7 0.69 0.71 $\pm$ 0.01 0.70 $\pm$ 0.00 6:02:45
MLP FashionMNIST 2 4 32 - 0.62 0.68 0.64 0.71 $\pm$ 0.02 0.68 $\pm$ 0.00 11:42
Table 4: Impact of KAN Parameter Adjustments on MNIST and FashionMNIST Performance
Model Dataset # Conv Layers # Filters # Width Grid Size With Feature Selection Training Time (hh:mm:ss)
maP maR maF1 Train Accuracy Test Accuracy
KAN MNIST 2 4 8 2 0.72 0.72 0.72 0.72 $\pm$ 0.02 0.72 $\pm$ 0.00 6:05:24
KAN MNIST 2 4 8 3 0.73 0.73 0.73 0.75 $\pm$ 0.02 0.73 $\pm$ 0.00 5:50:29
KAN MNIST 2 4 16 3 0.75 0.74 0.74 0.74 $\pm$ 0.02 0.75 $\pm$ 0.00 10:21:59
KAN FashionMNIST 2 4 8 2 0.70 0.71 0.70 0.72 $\pm$ 0.02 0.71 $\pm$ 0.00 5:11:51
KAN FashionMNIST 2 4 8 3 0.69 0.70 0.69 0.71 $\pm$ 0.01 0.70 $\pm$ 0.00 6:02:45
KAN FashionMNIST 2 4 16 3 0.71 0.72 0.71 0.73 $\pm$ 0.02 0.72 $\pm$ 0.00 11:41:30
5 Results
Our results reveal that KAN-DDQN outperforms MLP-DDQN in all cases on both MNIST and FashionMNIST datasets (see Fig. 3a and 3b) with 4 times fewer parameters and only $8\times 8$ sized images. Beyond top-1 accuracy, we report macro F1, macro AUC-PR, and calibration (ECE, NLL, Brier, with reliability diagrams) in Appendix LABEL:app:1.
5.1 Effect of Distribution Type in Feature Selection
Table 1 demonstrates significant variability in MLP-DDQN performance across different distributions for both the MNIST and FashionMNIST datasets. We conducted this experiment to choose the best distribution that can mimic the human-like feature selection process. The beta distribution consistently yields the highest train and test accuracies on MNIST (79% and 78%, respectively), prompting us to conduct the remaining experiments with this distribution. Conversely, the Gumbel Softmax, Dirichlet, multinomial, Cauchy, and Laplace distributions perform poorly, with test accuracies as low as 12%. Notably, the uniform distribution also shows competitive results, particularly on the MNIST dataset, suggesting it can serve as a viable alternative for feature selection.
5.2 Effect of Feature Selection in Performance
5.2.1 Performance With Feature Selection
Table 3 illustrates the performance of the KAN-DDQN and standard MLP-DDQN using feature selection across the MNIST and FashionMNIST datasets. In all cases, the hidden layer width was set to 32 neurons for the MLP and 8 for the KAN, reflecting the distinct architectures of each model. The KAN model consistently outperformed the MLP across all configurations, particularly on the MNIST dataset, achieving a test accuracy of 93% with a single convolutional layer, two filters, a width of 8, and a grid size of 3. For FashionMNIST, KAN similarly demonstrated superior test performance, reaching 83% accuracy under the same configuration, compared to MLP’s 74% accuracy. Across both datasets, KAN exhibited higher mean average precision (maP), recall (maR), and F1-score (maF1). For example, in the single-layer configuration on MNIST, KAN achieved scores of 0.92 for each metric, whereas MLP scored 0.84. The training time for the KAN model was generally longer due to its more intricate architecture, with the best-performing KAN model on MNIST requiring 4 hours and 35 minutes, compared to only 8 minutes for the MLP (see Table 3). Despite the increased computation, the accuracy gains suggest that KAN is efficient for applications where performance outweighs training time constraints. Notably, with the addition of a convolutional layer and increased filter size (e.g., 2 layers, 4 filters), KAN’s performance slightly declined, particularly on FashionMNIST, where test accuracy was 70%. This suggests a potential trade-off between model complexity and generalization, indicating that simpler KAN configurations, particularly those with a lower hidden layer width, may better retain discriminative power for smaller datasets.
5.2.2 Performance Without Feature Selection
Table 2 provides a comparative view of model performance without feature selection, highlighting the impact of excluding feature selection on both models’ performance. Without feature selection, both KAN and MLP demonstrate considerably poor performance. Here, the KAN model achieves a notable decrease in test accuracy and maF1 compared with the MLP model. Specifically, on MNIST, KAN reaches a test accuracy of 58% with maF1 of 0.56, surpassing the MLP model’s 52% test accuracy and 0.47 maF1. On FashionMNIST, the KAN model continues to outperform, achieving a 64% test accuracy and 0.63 maF1, compared with MLP’s 59% accuracy and 0.57 maF1.
5.3 Effect of KAN Parameters in Performance
Table 4 shows how KAN performance varies with parameter adjustments, focusing on width (number of neurons in the hidden layer) and grid size. Doubling the width of the feature maps from 8 to 16 yielded improvements in accuracy, especially on MNIST, where the test accuracy rose from 72% to 75%. However, this increase came at the expense of significantly longer training times, suggesting a trade-off between accuracy and computational efficiency. On FashionMNIST, similar parameter adjustments yielded more moderate gains. For example, increasing the grid size to 3 and the width to 16 raised test accuracy slightly from 71% to 72%. Moreover, altering the grid size from 2 to 3 produces varied outcomes: it preserves similar test accuracy in MNIST (0.72 to 0.73) but slightly decreases performance in FashionMNIST (0.71 to 0.70). This variation may stem from applying the B-spline in higher dimensions, warranting further exploration. This pattern suggests that KAN’s parameter sensitivity is dataset-dependent, with MNIST benefiting more from only width increases. Training time increased substantially with larger configurations. On MNIST, the KAN model with a grid size of 3 and a width of 16 took over 10 hours, while the smaller grid size of 2 and a width of 8 configurations was completed in 6 hours. Thus, reducing width and grid size may be advantageous for time-sensitive applications, especially where marginal accuracy improvements are not critical. So, larger hidden layer widths improve KAN’s accuracy but at the cost of increased training time. Balancing grid size and width for applications prioritizing efficiency may offer an optimal trade-off without sacrificing significant accuracy.
5.4 Cross-Dataset Performance and Scalability
A consolidated summary of results across MNIST, FashionMNIST, CIFAR-10, and CIFAR-100 is provided in Appendix 1010.1 and Table 5, highlighting consistent gains of KAN over MLP in accuracy, calibration, and scalability.
5.5 Ablation Studies
Ablation results are reported in Appendix 1010.2 and Table 6, confirming the critical role of instance-wise feature selection and differentiable gates in maintaining accuracy and calibration.
5.6 Complexity and Deployability Analysis
Model complexity, latency, and sparsity metrics are summarized in Appendix 1010.3 and Table 7, demonstrating that KAN remains deployable on edge devices despite moderate training overhead.
5.7 Interpretability of KAN-DDQN
In our KAN-DDQN architecture, several interpretability features improve the understanding of the model’s decision-making process, particularly after training on datasets.
5.7.1 Pruning
After training, we implement a pruning mechanism to streamline the KAN architecture by eliminating less important neurons (see Fig. 4a). The significance of a neuron is determined by its incoming and outgoing scores, defined as:
$I_{l,i}=\max_{k}(|\phi_{l-1,i,k}|_{1}),\quad O_{l,i}=\max_{j}(|\phi_{l+1,j,i}|_{1})$ (25)
A neuron is considered important if both its incoming score $I_{l,i}$ and outgoing score $O_{l,i}$ exceed a threshold $\theta=10^{-2}$. Unimportant neurons not meeting this criterion are pruned from the network, resulting in a more efficient model.
5.7.2 Visualization
To understand the importance of features, we utilize the plotting functionality of the KAN model. Fig. 4a visualizes the activation functions of the neurons, and Fig. 4c shows the training steps of the KAN agent, where the bottom layer indicates 64 input variables, followed by 8 hidden neurons, and then 10 outputs. The transparency of each activation function $\phi_{l,i,j}$ is set proportionally to $\tanh(\beta A_{l,i,j})$, where $\beta=30$. Smaller $\beta$ allows us to focus on more significant activations.
5.7.3 Symbolification
Our approach identifies symbolic forms within the KAN architecture to enhance interpretability. We sample network preactivations $x$ and post-activations $y$, fitting an affine transformation $y\approx cf(ax+b)+d$, where $a$, $b$ adjust inputs, and $c$, $d$ scale and shift outputs. This fitting involves grid search on $a$, $b$ and linear regression for $c$, $d$. Using a library of symbolic functions ($x,x^{2},x^{3},\exp(x),\log(x),\sqrt{x},\sin(x),|x|$), auto-symbolic regression replaces learned activations with interpretable formulas. The policy in KAN-DDQN is represented as $a_{i}=f_{i}(x)$ for each output $i$, with the overall action $a=\underset{i}{\mathrm{argmax}}\ a_{i}$. Fig. 4b displays the importance of features extracted from a symbolic formula, showing their frequency of occurrence and corresponding contributions to the model’s predictions. Fig. 5 illustrates the activation function of a middle neuron in the KAN’s 8-neuron hidden layer for two input neurons, with the x-axis showing input values and the y-axis indicating the activation output after the spline function.
6 Conclusion and Future Work
This study introduced a HITL feature selection framework using a DDQN architecture with KAN, achieving interpretable, per-example feature selection and improved performance across multiple metrics. Through simulated feedback and stochastic feature sampling from diverse distributions, the HITL-KAN-DDQN model dynamically selects feature subsets for each observation, focusing on the most relevant features and improving interpretability. Operating on low-resolution $8\times 8$ pixel images, our model consistently outperforms MLP-based DDQN models, achieving higher accuracy while using fewer neurons, specifically, an 8-neuron single hidden layer compared to the 32-neuron layer required by the MLP-DDQN. This compact design provided computational efficiency without compromising predictive power, confirming the effectiveness of the HITL-KAN-DDQN model for feature selection.
Future work will extend this HITL feature selection framework beyond low-resolution vision benchmarks to include standard-resolution images and non-vision modalities such as tabular and multimodal datasets. This will enable rigorous evaluation of generalization across diverse, real-world scenarios and characterize the benefits and limitations of instance-wise selection under higher-dimensional and noisier conditions. We also aim to improve simulated feedback mechanisms to better align with expert annotations, further strengthening interpretability and fairness. To address the computational demands of KAN models, future implementations will incorporate more efficient variants such as FastKAN [10] and PowerMLP [12], leveraging their lightweight architectures to reduce training time while maintaining interpretability and performance. Another critical direction is to replace simulated feedback with actual human feedback from domain experts in future iterations. By involving real experts during the interactive feature selection process, we can increase the practical relevance, interpretability, and fairness of the selected feature subsets. Integrating expert-driven annotations will also facilitate comparative analysis between simulated and authentic feedback systems, enabling refined alignment strategies. From a reinforcement learning perspective, future research will explore active learning-based query strategies that prioritize feedback acquisition based on model uncertainty and data point importance. Moreover, we intend to experiment with a broader range of reinforcement learning algorithms beyond DDQN, including DQN, REINFORCE, proximal policy optimization (PPO), advantage actor-critic (A2C), and soft actor-critic (SAC), to evaluate their suitability for adaptive, per-example feature selection in human-in-the-loop settings.